OpenStudy (elite):

Which regular polygon would have each of its interior angles measure 120°?

OpenStudy (studygurl14):

Interior Angle formula: \(\Large\frac{180(n-2)}{n}\) where n = number of sides

OpenStudy (studygurl14):

So... \(\Large 120 = \frac{180(n-2)}{n}\) solve for n